# Hypergraph Representations of scRNA-seq Data for Improved Clustering with Random Walks

### v2

@@ I Introduction
While these methods employed unipartite networks to aid analysis, it is often overlooked that unipartite networks are often not the most natural or information-preserving representation of the complex system we are interested in studying since many complex systems involve interactions from more than one type of entities. Examples include protein-protein interaction (PPI) networks where the nodes are proteins and edges could be physical contacts among proteins or functional interactions in biological pathways, gene regulatory networks (GRNs) where nodes represent genes and a directed edge exists if one gene regulates the expression of another gene, co-expression networks where the nodes are genes, cells, or tissues and the weighted edges indicate similarity in expression profile between node pairs.
An example of a network that captures higher-order interactions between different entity types is the bipartite network, where nodes are partitioned by two separate groups, and edges only connect nodes from different groups. Bipartite networks are particularly suitable for modeling systems where two types of entities interact, such as authors and papers in a collaboration network [18], recommendation systems [19] where nodes represent users and the recommending items, and specifically in biology, plant-pollinators and predator-prey interaction networks [20] in community ecology studies, microbe-host interaction networks [21] in microbiology, drug–protein interaction networks [22] , and transcription factor-gene networks in genomics.
However, instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14]. For example, in the bipartite representation of a drug interaction network, the two classes of nodes represent drugs and protein targets respectively. A drug node and a protein node are connected if the drug targets the protein. In its unipartite transformed network, a drug-drug interaction network could be constructed based on shared targets. Similarly, in the bipartite protein complex network, proteins and biological functions are two types of nodes that are connected if the proteins are involved in the biological function. Yet, in the unipartite representation of the protein complex network, the clique graph of the original bipartite graph is constructed where the only class of nodes are the proteins and the protein nodes are connected if they are in the same complex that performs a biological function. Unipartite representations fail to capture the biological functions associated with groups of proteins. While unipartite projections can reveal relationships between nodes of the same type in a bipartite network, they often obscure the specificity of these interactions.
@@
Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data. Hypergraphs offer a neat solution by representing multi-way relationships in scRNA-seq directly, without the need for further data projection. Moreover, each bipartite graph can also be represented as a hypergraph, by modelling one set of entities as nodes and the other as hyperedges, offering a generalized representation of high-order interactions. In this work, we propose a hypergraph-based approach to analyze scRNA-seq data to overcome the limitations of unipartite projections.
@@ II Methods
In scRNA-seq analysis, clustering cells based on co-expression networks projected from count data is a common approach for cell type identification. However, as discussed in III.3, the correlation inflation problem associated with co-expression networks caused by the induced sparsity in scRNA-seq data, as well as information loss from projecting a hypergraph to a unipartite simple graph, motivated us to look beyond co-expression networks for scRNA-seq analysis.
We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$. In this framework, we propose a hypergraph random walk-based clustering method that models the relationships between cells and genes in the cell-gene expression abundance count data. Specifically, each cell is represented as a node, and each gene is represented as a hyperedge. Each hyperedge captures the expression profile of a gene across all cells, allowing us to apply hypergraph based analysis tools without enduring the loss of information caused by projecting the scRNA-seq data into a simple unipartite graph, e.g. the co-expression networks. We design random walk algorithms that capture cell-cell interactions by sampling the cell-to-gene and gene-to-cell paths in the conceptualized hypergraph. The random walk sampling probability distribution is designed to account for the specific cell-gene expressions via the edge-dependent vertex selection probability and the vertex-dependent edge selection probability. The sampled paths embed information regarding the amplified expression level of a gene in a cell, both in comparison to the expression levels of other genes in that cell and to the expression levels of the same gene in other cells. As a step further, CoMem-DIPHW (the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk), the second hypergraph-based random walk algorithm proposed for clustering, incorporates a memory mechanism that takes into account both the conceptualized hypergraph from the scRNA-seq expression data, and the two projected unipartite co-expression networks, i.e. the cell and gene co-expression network, for cell-cell interaction sampling. The sampled cell-cell interaction paths would then be fed to a small language model (Word2Vec) to generate efficient numerical representations (embeddings) of the cells. These embeddings would then be used to cluster the cells by applying a clustering algorithm such as UMAP or t-SNE to group the cells based on their similarity in the embedded space.
We emphasize that this hypergraph conceptualization of the scRNA-seq data does not require further transformation on the data and consequently would not result in any information loss or computation cost from converting the scRNA-seq into co-expression networks. We remark that a matrix is a 2-dimensional array structured for organizing information related to two variables. Thus a matrix could naturally be conceptualized as a hypergraph by considering the column variables as nodes and the row variables as edges and vice versa.
@@ II.1 Hypergraph representation
  • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed.
The incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ is constructed such that $\mathbf{I}_{\mathcal{H}}(v,e)$ is given by the expression level, i.e. abundance counts, of gene $e$ in cell $v$.
II.2 Bipartite representation
@@
II.3 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
We introduced a novel hypergraph random walk approach, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), highlighting the consideration of both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. In a hypergraph random walk process, the walker typically alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$. In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the previously visited node, i.e., $P_{V|E}(v,e)=P_{E}(e)$ for all $v\in V$. As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process. The modified random walk process is described as follows: at time $t$, assuming the random walker is at vertex $v_{t}$, the walker proceeds by iteratively selecting the next hyperedge and the subsequent vertex within that hyperedge to transition to, based on the following probabilities:
@@ II.3 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
    In the context of scRNA-seq cell clustering, this cell-dependent gene selection probability depends on the relative expression level of a gene in a cell compared to the expression of other genes in the same cell.
@@
    Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is defined as:
@@
    In the context of scRNA-seq cell clustering, the transition from a gene to a cell is based on the relative expression level of a gene in a cell, amplified by the preference exponent for faster clustering convergence, and normalized by the same gene’s expression in other cells.
@@
    Node-to-Node Transition Probability: Marginalizing over all possible hyperedges the walker can traverse gives the unipartite node-to-node transition probability of DIPHW. This probability is used for sampling the random walk node paths which capture the structural and relational information of the original hypergraph and are subsequently used as inputs to a neural network for cell embedding computation.
@@
    For efficient computation, the $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
@@
    The vectorization circumvents the need for explicit iteration and reduces the computation time. Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ is the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$. Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization, with dimensions $|E|\times|V|$. Each entry in this matrix is obtained by raising the corresponding edge-dependent vertex weight to the power determined by the preference exponent $\alpha$, expressed as $W_{V|E}=(\mathbf{I}_{\mathcal{H}}^{\alpha})$, where $w_{ve}=[\mathbf{I}_{\mathcal{H}}(e,v)]^{\alpha}$. Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal normalizing matrices used to normalize the distribution of the transition from vertex-to-edge and edge-to-vertex respective so that the probability distribution sums to 1.
II.4 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
In this section, we introduce the co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW) designed to integrate scRNA-seq expression data and its co-expression networks for improved cell embedding computation. CoMem-DIPHW builds upon the Dual-Importance Preference Hypergraph Walk (DIPHW) and incorporates a memory mechanism to consider previously visited nodes and edges during transitions.
Hypergraph Representation and the unipartite network projections of the scRNA-seq data
We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without the need for any additional computation. Moreover, we can construct two correlation networks as derivatives of the matrix. In each correlation network, the nodes correspond to one of the two variables, and the edges represent the correlations between the node pairs of the corresponding variable. These two unipartite correlation network projections embed global level information that summarizes the association between node pairs through their correlation.
@@
Using scRNA-seq data, we construct two co-expression networks: the gene-co-expression network (GCN) $G_{E}$ and the cell-co-expression network (CCN) $G_{V}$. In the GCN, genes are represented as nodes, and edges correspond to correlation coefficients between gene pairs, indicating the similarity between their expression profiles. Similarly, in the CCN, cells are represented as nodes, and edges correspond to correlation coefficients between cell pairs. These two co-expression networks are unipartite projection networks of the original scRNA-seq expression data, with edge weights providing a summary statistic of pairwise similarity in expression profiles. Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it is a non-parametric measure that captures non-linearity compared to Pearson’s correlation coefficient. The cell and gene co-expression networks respectively are the node and edge similarity networks derived from the hypergraph. These two unipartite correlation network projections embed global level information that summarizes the association between cell and gene pairs through their correlation.
@@
Our proposed random walk process CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges. This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited. The node-node and edge-edge dependencies are incorporated through the node-node and edge-edge similarity networks, which are the two correlation networks derived from the original matrix. The memory incorporation allows both the local and global structures in the hypergraph to be captured by the walk. Moreover, the memory mechanism constrains the hypergraph walker to focus on network(cell-gene) interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
Capturing Global and Local Geometry
In our proposed CoMem-DIP-HW random walk, we leverage global information from unipartite projections—the gene-co-expression network $G_{E}$ and the cell-co-expression network $G_{V}$—along with local information from the hypergraph structure in the cell-gene scRNA-seq expression data to enhance clustering performance. The co-expression networks serve as node and edge similarity matrices, facilitating the memory-incorporated hypergraph random walk. For cell clustering and cell type identification in scRNA-seq data, cells are modeled as nodes and genes as hyperedges. The hypergraph represents the expression profiles of cells across genes, where each hyperedge (gene) connects to nodes (cells) in which it is expressed. The weight of connection in a hyperedge to each node (cell) is determined by the gene’s expression level in that cell. The cell co-expression network $G_{V}$ functions as the node similarity matrix, with $G_{V_{i,j}}$ representing the similarity between the expression profiles of cell pair $i$ and $j$. Similarly, the gene co-expression network $G_{E}$ serves as the edge similarity matrix, representing the similarity in expression profile between pairs of genes across different cells. Each entry in these matrices aggregates information from the expression profiles of pairs of cells or genes into pairwise similarities, providing a global perspective through unipartite network projections. However, specific expression levels of genes in individual cells, representing local information, are lost in these projected networks. Therefore, our proposed hypergraph random walk methods aim to integrate global information while preserving local information, to achieve more accurate cell type identification.
@@
Finally, the unipartite node-to-node transition probability of the multilayer network random walk process is:
$P_{V_{t}\rightarrow V_{t+1}(v,w)}$
$\propto$ $\sum_{e_{1},e_{2}\in E}P_{V|E}(v|e_{1})P_{E_{2}|V,E_{1}}(e_{2}|v,e_{1})P_{V_{t+1}|E,V_{t}}(w|e_{2},v)$
@@
Further, for CoMem-DIPHW, we incorporate DIPHW proposed in section II.3 to CoMem walk II.4. We define
  • The probability of selecting vertex $v$ in edge $e$ is:
    $P_{V|E}(e\rightarrow v)=\frac{\gamma_{e}(v)^{\alpha}}{\sum_{v^{\prime}\in e}\gamma_{e}(v^{\prime})^{\alpha}}$
  • The probability of selecting edge $e$ given vertex $v$ is:
    $P_{E|V}(v\rightarrow e)=\frac{\omega(e)\gamma_{e}(v)}{\sum_{e^{\prime}\in E(v)}\omega(e^{\prime})\gamma_{e^{\prime}}(v)}$
    The other probability distributions are the same as defined in section II.4
II.5 Complexity Analysis
DIPHW
$P_{V|V}=D_{E|V}^{-1}W_{E|V}D_{V|E}^{-1}W_{V|E}$
CoMem-DIPHW
@@ III.1 Baseline Methods
To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them to several widely used and established cell clustering methods, including five embedding-based methods (PCA [43], Node2Vec [44], EDVW [41], UMAP, and t-SNE) and four unipartite-graph-based community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). These methods are adopted by popular scRNA-seq analysis packages such as Seurat [15], WGCNA [14], and Scanpy [17]. We provide a summary of these methods in Table 2. For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy. This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps in the package implementations. The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
@@ III.2 Data and Preprocessing
The data used in this study consists of both simulated and publicly available scRNA-seq datasets. The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56]. We have provided a detailed README file to guide users through the process of running the code and reproducing our results. For preprocessing, the raw single-cell RNA-seq data were first normalized to equalize total counts per cell to account for differences in sequencing depth among cells. After normalization, genes and cells with zero expression were removed. For empirical analysis, we retained only the top n most variable genes to focus on the most relevant signals. For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].

### v3

@@ I Introduction
However, traditional methods represent the biology systems using unipartite networks, which only captures relationships (edges) among one type of entity (nodes), are often not the most natural or information-preserving representation of the complex systems since many involve interactions between multiple types of entities. Instead of modelling these systems using network models that could capture higher order interactions, most often, a unipartite network projection is chosen such that the system could be analyzed with the abundant network analysis methods developed based on unipartite networks [23, 14]. For example, in protein-protein interaction (PPI) networks, unipartite representations capture relationships between proteins but fail to capture the biological functions associated with groups of proteins. Similarly, unipartite projections of gene regulatory networks connect genes based on regulatory relationships but do not capture relationships among different entities such as transcription factors or their binding sites.
@@
Given the issues with unipartite projections, it is crucial to explore alternative network representations that can accurately reflect the scRNA-seq data. Hypergraphs offer a neat solution by representing multiway relationships in scRNA-seq directly, without the need for further data projection.
@@ II Methods
We conceptualize scRNA-seq data as a hypergraph $\mathcal{H}$ and design two hypergraph random walk algorithms that capture cell-cell relationships by sampling the cell-to-gene and gene-to-cell interactions in the conceptualized hypergraph. We emphasize that this hypergraph representation of the scRNA-seq data does not require further transformation of the data and consequently would not result in any information loss or additional computational cost.
@@ II.1 Hypergraph representation
  • $E$ denotes the hyperedge set representing genes, with each hyperedge $e\in E$ connecting to all cells $\in V$ in which the gene $e$ is actively expressed. The weight of connection of a hyperedge (gene) to each node (cell) is determined by the gene’s expression level in that cell.
Thus, each hyperedge captures the expression profile of a gene across all cells, preventing any loss of information. The incidence matrix $\mathbf{I}_{\mathcal{H}}$ of the hypergraph $\mathcal{H}$ is constructed such that $\mathbf{I}_{\mathcal{H}}(v,e)$ is given by the expression level, i.e. abundance counts, of gene $e$ in cell $v$.
@@
II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
Building on this hypergraph representation, we introduced a novel hypergraph random walk approach for cell embedding computation, the Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW), which accounts for both the relative importance of edges to nodes and nodes to edges, or importance of genes relative to cells and cell to genes, along with a preference exponent for faster clustering. In a hypergraph random walk process, the walker alternates between transitioning from a hyperedge to a node and from a node to a hyperedge, following the node-to-edge transition probabilities $P_{E|V}(e,v)$ and edge-to-node transition probabilities $P_{V|E}(v,e)$. In the hypergraph random walk with edge-dependent vertex weight (EDVW) by Chitra and Raphael (2019) [41], the edge-to-node transition probability is edge-dependent, but the node-to-edge transition probability does not consider the importance of the edge to the node. As a natural extension to EDVW, we defined the node-to-edge transition probability to be vertex-dependent in DIPHW, and used a preference exponent to accelerate the clustering process. The modified random walk process is described as follows:
@@ II.2 Method 1: Dual-Importance Preference (DIP) Hypergraph Walk (DIPHW)
    In scRNA-seq cell clustering, this represents the probability of selecting a gene based on its expression level relative to other genes in the same cell.
@@
    Upon the walker’s arrival at hyperedge $e_{t+1}$, the probability of moving to vertex $v_{t+1}\in e_{t+1}$ is given by:
@@
    In scRNA-seq cell clustering, the transition from a gene to a cell is based on the relative expression level of a gene in a cell, amplified by a preference exponent to improve clustering convergence.
@@
    Node-to-Node Transition Probability:
    The unipartite node-to-node transition probability of DIPHW is obtained by summing the product of the node-to-edge and edge-to-node transition probabilities over all hyperedges. This probability is used to sample random walk node paths, which serve as inputs to a neural network for cell embedding computation.
@@
    The $|V|\times|V|$ node-to-node transition probability matrix $P_{V|V}$ could be written in the vectorized form as,
@@
    Vectorization eliminates explicit iteration and reduces the computation time. Here, $W_{E|V}=\mathbf{I}_{\mathcal{H}}^{T}W_{E}$ is the $|V|\times|E|$ node-to-edge transition weight matrix where column $e$ of $W_{E|V}$ contains the vertex weights in hyperedge $e$ multiplied by the weight of hyperedge $e$. Similarly, $W_{V|E}$ represents the edge-to-node transition weight matrix before normalization with dimensions $|E|\times|V|$, given by $W_{V|E}=\mathbf{I}_{\mathcal{H}}^{\alpha},\quad w_{ve}=\left[\mathbf{I}_{\mathcal{H}}(e,v)\right]^{\alpha}$. Finally, $D_{E|V}$ and $D_{V|E}$ are diagonal matrices that normalize vertex-to-edge and edge-to-vertex transitions to ensure the probability distribution sums to 1.
II.3 Method 2: co-expression and Memory-Integrated Dual-Importance Preference Hypergraph Walk (CoMem-DIPHW)
We now introduce CoMem-DIPHW, an extension of DIPHW that incorporates co-expression networks and single-cell transcriptomic profiles. The memory mechanism of CoMem-DIPHW leverages cell and gene co-expression networks to account for previously visited nodes and edges during transitions to capture both local expression relationships and global co-expression patterns.
Hypergraph Representation of the scRNA-seq data
We remark that a matrix can be directly interpreted as the incidence matrix of a hypergraph to represent the relationship between two variables, without additional computation.
@@
Using scRNA-seq data, we construct two co-expression networks: the cell-co-expression network (CCN) $G_{V}$ and the gene-co-expression network (GCN) $G_{E}$. In the CCN, cells are nodes, and edges correspond to correlation coefficients between cell pairs, indicating the similarity between their expression profiles. Similarly, in the GCN, genes are nodes and edges correspond to correlation coefficients between gene pairs. Spearman’s rank coefficient was chosen to measure pairwise similarity of expression profiles as it captures non-linearity better than Pearson’s correlation. These unipartite networks serve as node and edge similarity matrices, embedding global-level information that summarizes associations between cell and gene pairs.
@@
CoMem-DIPHW integrates a memory component by considering the similarity between consecutively visited nodes and edges. This memory mechanism adjusts the transition probabilities such that the node-to-edge transition probability depends on the previous edge visited, and the edge-to-node transition probability depends on the previous node visited. The node-node and edge-edge dependencies are incorporated through the node and edge similarity networks, which are the two correlation networks derived from the transcriptomic expression data. The memory mechanism constrains the hypergraph walker to focus on cell-gene interactions among similar nodes(cells) and edges(genes) to better uncover modularity within the data.
While the co-expression networks provide global perspectives through unipartite projections, specific expression levels of genes in individual cells (local information) are preserved in the hypergraph structure. CoMem-DIPHW integrates both global and local information for more accurate cell type identification.
@@
Finally, the unipartite node-to-node transition probability of the hypergraph random walk process is:
$P_{V_{t}\rightarrow V_{t+1}(v_{1},v_{2})}$
$\propto$ $\sum_{e_{1},e_{2}\in E}P_{V|E}(v_{1}|e_{1})P_{E_{2}|V,E_{1}}(e_{2}|v_{1},e_{1})P_{V_{t+1}|E,V_{t}}(v_{2}|e_{2},v_{1})$
@@
For CoMem-DIPHW, we integrate DIPHW into the CoMem walk framework from section II.3. Specifically, the vertex selection probability $P_{V|E}(e\rightarrow v)$ and the edge selection probability $P_{E|V}(v\rightarrow e)$ are both defined as in DIPHW (section II.2), while all other probability distributions remain as specified in section II.3.
@@ III.1 Baseline Methods
To evaluate the performance of DIPHW and CoMem-DIPHW, we compared them to several widely used and established cell clustering methods, including five embedding-based methods (PCA [43], Node2Vec [44], EDVW [41], UMAP, and t-SNE) and four unipartite-graph-based community detection methods (Louvain, Greedy Modularity, Infomap, and Leiden). These methods are adopted by popular scRNA-seq analysis packages such as Seurat [15], WGCNA [14], and Scanpy [17]. We provide a summary of these methods in Table 2. For a fair comparison under consistent conditions, we focus on the core clustering algorithms rather than specific implementations in packages such as Seurat and Scanpy. This allows us to evaluate the clustering capabilities of each method independently of any additional preprocessing or optimization steps. The hyperparameters used for each experiment are detailed in Table S1 in the supplementary materials Supplementary Materials.
@@ III.2 Data and Preprocessing
The data used in this study consists of both simulated and publicly available scRNA-seq datasets. The simulated data were generated using our proposed simulation algorithm described in detail in the Methods section, with code available on GitHub [56]. For preprocessing, we first normalized the raw single-cell RNA-seq data to account for differences in sequencing depth. Genes and cells with zero expression were then removed, and only the top n most variable genes were retained for analysis. For empirical analysis, we used four publicly available scRNA-seq datasets: human brain [57], human pancreas [58], mouse brain [59], and mouse pancreas [60], and a curated benchmarking dataset with ground-truth cell line labels [61].